Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day04


Crew 97 Log Book for January 4, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status: 
Overall ok. A few crew-members have minor conditions of a cold. 

Time Departed/Returned from EVA: 
EVA #1: 10:30 / 11:00. Three Crew Members 
EVA #2: 15:10 / 16:03. Two Crew Members 

Brief Narrative of Field Mission Results: 
Crew successfully collected pebble rock and fine sand samples for use in creating a slow sand filter for Judah's Sedimentary Water Filter project. 

EVA Data/Interpretations: 
EVA #1: Three Crew Members. Crew traveled in the vicinity of the HAB to the North and collected pebble sized rock samples for use in creating a slow sand filter. 

EVA #2: Two Crew Members. Crew traveled up to Skyline Rim to find sandstones which will be broken up into sand particles for use in creating a slow sand filter. 

Engineering/Hab Maintenance: 
EVA radios are functioning. 
Greywater line to toilet is now currently functioning. 
The water line to the indoor HAB tank is not yet functioning.  
Maintenance is being conducted on the ATVs. 

Report Transmission Schedule: 
CMDR Check-in, ENGR, CMDR Report, Photos, 2 EVA Reports, Science Report 

Plans for Tomorrow: 
- 07:30 - Breakfast (Two Crew Members) 
- 8:00 - Crew exercise 
- Engineering checkups 
- EVA #3 
- 11:30 - Lunch 
- Engineering - work on ATVs, and water pumps 
- EVA #4 
- Preparation of crew experiments 
- Engineering checks 
- 17:30 - Dinner 
- Reports by 20:00 
- Food Study Questionnaire 
- Showers: Two Crew Members 

Inventory: 
n/a 

Support Requested: 
DG plans bring a new hose for the potable water pump. 

Miscellaneous: 
n/a 


Commander's Report

Judah Epstein Reporting

Today the crew awoke to life on Mars. The crew is becoming accustomed to our routine of 7:30am breakfast followed by morning exercise. A crew member then performs engineering checks on the HAB. Afterwards, the crew prepared for and began the first EVA of Crew 97. Three crew members ventured outside in the spacesuits to collect pebble rock samples for use in creating a slow sand filter. 


After lunch, the crew was busy with various tasks such as engineering repairs on the ATVs and successfully repairing the greywater flush line to the toilet. Additional preparation was conducted towards scientific experiments such as physiological preparation and setup of the rover. Two crew members further explored Mars through EVA #2 of Crew 97. The team trekked up to Skyline Rim to collect sandstone rocks. The rocks were carried back to the HAB and will be broken into sand particles for use as the filter media of Judah's slow sand filter experiment. This slow sand filter is currently being created and will be compared with the results of two other already created slow sand filters with an established schmutzdecke biological layer (Martian simulant sand sediment and commercially available sand sediment); with grey water influent. 


A crew member has received the Food Study Questionnaire and is coordinating the crew's adherence to the protocols of the Food Study. 


Overall the crew is in good spirits and is exciting about the Simulation and scientific projects.



Science Report

Lucinda Land Reporting

Unpacked the pelican case that was delivered a few days ago to the Hab. Too many other responsibilities to do such as getting familiar with the Hab systems and kitchen rotation duties before I was able to get to it. 


Before unpacking the rover I wanted to take pictures of what it looked like packed so that I can fit everything back into its box for when I return ship it. I created a map of what contents were where. I then removed piece by piece and took inventory of all the parts and compared them to the packing list. The USB hub was the only item missing. There was also another tool missing, but after some communication with our team back at NASA, I was told where I might find it. I have not as of yet confirmed the location of that part. 


It looks like I have everything I need to get started. Tomorrow I will start the rover named, "Sandstorm" up. I want to do some preliminary testing before taking her out I want to make sure that she at least starts up successfully. She has been here twice before and was unsuccessfully utilized for unknown reasons. I am confident that she will be successful this time around. We shall see tomorrow.


Engineering Report

Nathan Wong Reporting

Generator/Electricity: 


Kitty(Diesel Generator) (Used? Yes/No): Yes 


Honey (Gas Generator) (Used? Yes/No): No 


Battery Charge Level: 


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): All systems operational 



Fuel Status: 


Diesel (Full/0.75/0.5/0.25/Empty): 0.75 


Propane Tank (Full/0.75/0.5/0.25/Empty): 0.64 


Gasoline Tank (Full/0.75/0.5/0.25/Empty): 4 Gas cans full 


Notes/Comments: 



Oil: 


Kitty Oil Quantity (# of Quarts in storage): 3 


Honey Oil Quantity (# of Quarts in storage): 2 


ATV Oil Quantity (# of Quarts in storage): 2 


Car Oil Quantity (# of Quarts in storage): none 



Water Status: 


Outside Potable Water Tank Level (inches from bottom): 25% 


Trailer Potable Water Tank Level (inches from bottom): empty 


Hab Potable Water Tank Level (inches from bottom): 80% 


Potable Water Meter Reading: 40323.6 


In to GreenHab Meter Reading: 


Notes/Comments: 


-Grey water line is now functional. 

-Pump to Hab tank is not functioning 

-Septic and grey water tanks have been emptied 



GreenHab: 


Condition of plants in Tank 1: N/A 


Condition of plants in Tank 2: N/A 


Condition of plants in Tank 3: N/A 


Greenhab Notes/Comments: 


Plants not at Hab yet. 



Observatory: 


Telescope Used? (Yes/No): No 


Observatory Notes/Comments: 


Observatory not used 



Transportation: 


Hab Car (NOT RENTAL) 

Used (yes/no): No 

Oil Checked (yes/no): No 

Tire Status: good 

Odometer at end of day: 123085 

Notes/Comments on Hab Car: 


ATV 1 (Opportunity): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 2 (Spirit): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 3 (Viking I): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 4 (Viking II): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


Viking II was the only operational ATV. Spirit and Opportunity were close to starting, and Viking I was not close to starting. This is the same as yesterday. I will keep working on them. 



Heating and Ventilation: 

Thermostat setting upstairs (Farenheit): 67 

Actual temperature upstairs (Farenheit): 75 


Thermostat setting downstairs (Farenheit): 63 

Actual temperature downstairs (Farenheit): 57 



Computers/Networking Infrastructure: 


Notes/comments on internet/computers?: 

Everything is working well 



General Engineering Notes/Comments: 


Everything is working well except the water line to the indoor Hab tank.


EVA 1 Report

Judah Epstein Reporting

Time: 10:30 / 11:00 


Crew Members: Judah, Nathan, Jim 


Site Location: In the vicinity of the HAB, to the North. Sample collection site was at N 38~24.500'. W 110~47.444. Elevation 4529 ft. 


Transit Mode: Foot 


Objectives: Collect pebble sized rock samples for use in creating a slow sand filter. 


Results: Rock samples were successfully collected and will be used as the filter base support. 


Lessons Learned: It is very difficult to see if the helmet is fogging. More difficult than expected to move around and communicate with crew-members while in EVA suit.



EVA 2 Report

Judah Epstein Reporting

Time: 15:10 / 16:03 


Crew Members: 2 


Site Location: Skyline Rim to the West of the HAB. Sample collection site was at N 38~24.442'. W 110~47.657'. Elevation 4611 ft. 


Transit Mode: Foot 


Objectives: Collect sandstones which will be broken up into sand particles for use in creating a slow sand filter. 


Results: Sandstones were successfully collected and will be used as the filter media. 


Lessons Learned: The crew was looking for fine sand, but did not find any soil of sand size. Crew-members improvised by collecting sandstone rocks/boulders to carry back to HAB to smash into sand particles.



Cook's Report

Judah Epstein Reporting

1) Was today 

- a cooking day? yes 

- a day of celebration involving special foods? No 


2) List the foods served at today's main meal (usually the evening meal), giving full names of each. 

Spaghetti with tomato/vegetable sauce made from tomato powder, seasonings and assorted rehydrated vegetables such as green beans, tomatoes, spinach, peas. 

Bear Creek Santa Fe Chipotle soup. 


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded if inedible) 

About 1 serving of extra spaghetti. 

About 1.5 servings of extra soup. 


4) Recipes for anything more complicated than following package directions exactly, or rehydrating. 

Tomato sauce with vegetables – recipe not reported 


5) Comments/questions 

- any shortages of food supplies or equipment? (includes "cannot find..." ) no 

- any problems with supplied food? (e.g. spoilage, insect or rodent damage, off flavors, difficulty rehydrating) no 

- any problems with kitchen equipment, including cleaning supplies? no 

- any cooking questions you need help with? no - Tonya assisted me 

- any other thoughts or suggestions? No